Simplify the following Boolean expressions using Boolean algebra. Show the simplification steps. a) ?(?̅? + ??̅) + ?(?? + ??̅) b) (? + ?)(?? + ??̅) + ?? + C
Answer #1

Part (a)

B(A'C + AC') + A(BC + BC')

= A'BC + ABC' + ABC + ABC' [By expanding]

= A'BC + ABC' + ABC [as ABC' + ABC' = ABC']

Since ABC + ABC = ABC so, we can substitute ABC by ABC + ABC

= A'BC + ABC' + ABC + ABC

= ABC' + ABC + ABC + A'BC

= AB(C + C') + BC(A + A')

= AB + BC [as C + C' = 1 and A + A' =1]

= B(A + C)

Part (b)

(A + C)(AD + AD') + AC + C

= (A + C)A(D + D') + AC + C

= (A + C)A + AC + C [as D + D' = 1]

= A + AC + AC + C

= A + AC + C [as AC + AC = AC]

= A(1 + C) + C

= A.1 + C [as 1 + C = 1]

= A + C

Hope this helps.